# Error-Driven Uncertainty Aware Training

### v1

@@ Abstract
Neural networks are often overconfident about their predictions, which undermines their reliability and trustworthiness. In this work, we present a novel technique, named Error-Driven Uncertainty Aware Training (EUAT), which aims to enhance the ability of neural models to estimate their uncertainty correctly, namely to be highly uncertain when they output inaccurate predictions and low uncertain when their output is accurate. The EUAT approach operates during the model’s training phase by selectively employing two loss functions depending on whether the training examples are correctly or incorrectly predicted by the model. This allows for pursuing the twofold goal of i) minimizing model uncertainty for correctly predicted inputs and ii) maximizing uncertainty for mispredicted inputs, while preserving the model’s misprediction rate. We evaluate EUAT using diverse neural models and datasets in the image recognition domains considering both non-adversarial and adversarial settings. The results show that EUAT outperforms existing approaches for uncertainty estimation (including other uncertainty-aware training techniques, calibration, ensembles, and DEUP) by providing uncertainty estimates that not only have higher quality when evaluated via statistical metrics (e.g., correlation with residuals) but also when employed to build binary classifiers that decide whether the model’s output can be trusted or not and under distributional data shifts.
@@ 2 Related Work
One important foundation of these works lies in the computation of uncertainty. Uncertainty in DNNs plays a pivotal role in quantifying the reliability and robustness of their predictions. There are two main types of uncertainty: epistemic uncertainty, associated with the model’s lack of knowledge or data, and aleatoric uncertainty, linked to the inherent randomness and unpredictability within the training data. Various metrics are employed to measure these types of uncertainty such as predictive entropy (PE) and mutual information (MI) [1]. However, quantifying uncertainty with DNNs is a challenging task. Bayesian methods [34, 28] can directly provide an estimate of the uncertainty by parameterizing the parameters of the network with distributions. However, training BNNs usually comes with a prohibitive cost. Thus, several approximations have been developed. Monte Carlo (MC) dropout [10], which is one of the most popular techniques for uncertainty quantification on DNNs [1, 4], proposed as a Bayesian approximation to estimate the uncertainty by sampling multiple dropout masks during training and inference phases and aggregating the predictions, providing a probabilistic estimate that allows the quantification of uncertainty. One can also approximate and estimate model uncertainty by computing the variance of the output predictions [5, 9, 25]. Additionally, Markov Chain Monte Carlo (MCMC) methods [12] offer another avenue for drawing the posterior distribution (albeit, those require a long time to converge to the final distribution [1, 34]). Further, Variational Inference (VI) is a technique used to estimate the uncertainty of BNNs by approximating the posterior distribution over the model’s weights, which are treated as random variables with associated probability distributions. Training the network to approximate these distributions enables the capture of inherent uncertainty in the model’s predictions. Moreover, we can also resort to models that output an uncertainty estimation, such as Deep Gaussian Processes [6], or use Laplace approximations for uncertainty  [29], or deep ensemble methods that offer yet another approach to estimate model uncertainty [24, 1], by aggregating the outputs of each learner in the ensemble and computing relevant metrics such as the entropy or MI. At last, to account for the computation overhead of the previous methods (resulting, for example, from training multiple models to the ensemble or performing several forward passes using MC dropout), DEUP [23] trains a new simple estimator to predict the uncertainty of the base model exploiting its error/loss, and DUN [2] leverages the outputs of different layers of a DNN to measure the uncertainty.
@@
Model calibration becomes especially critical in tasks where estimated uncertainties play a fundamental role in risk assessment. Thus, a first set of works aims at calibrating a fully trained model by applying a post-processing phase to align the output probabilities with the true likelihood of the predicted outcomes of events occurring [14]. DNNs are not inherently calibrated, and several techniques, such as Platt Scaling [39], Isotonic Regression [45], Temperature Scaling [14], or Beta Calibration [22], can be employed to fine-tune the probabilities outputted by the model ensuring a more accurate alignment with true outcome frequencies. Moreover, Krishnan et al. [20] introduced an accuracy versus uncertainty calibration (AvUC) loss function devised to obtain well-calibrated uncertainties while simultaneously preserving or enhancing model accuracy, and they extended their approach by proposing a post-hoc calibration phase that extends Temperature Scaling using AvUC. In addition, Karandikar et al. [18] proposed an extension of ECE and AvUC by developing a soft version of the binning operation underlying these calibration-error estimators, and also extended their approach for post-processing calibration by optimizing the temperature parameter in the temperature scaling method based on the soft calibration error. Complementary, Gupta et al. [15] presented a binning-free calibration approach. However, these calibration methods can be very sensitive to both the model and the validation set [27] and perform suboptimally when faced with shifts in data distribution [37].
@@
One fundamental aspect of training DNNs is the choice of a loss function. Although CE stands out as a common choice for addressing classification problems, it tends to increase the over-confidence of the resulting model [14, 32]. Thus, since accounting for predictive uncertainty during training improves model calibration [20], several loss functions have been developed to extend existing ones, aiming to incorporate the model’s uncertainty. These learning-based methods such as focal loss [26] or label smoothing [26] introduce an additional term addressing model uncertainty. These terms are typically balanced using corresponding weights introducing as a consequence additional hyper-parameters. Similarly, Shamsi et al. [41] proposed two loss functions that extend the CE by adding a new term to address the model’s uncertainty that can be determined through the PE or the ECE. Additionally, Einbinder et al [8] introduced an uncertainty-aware conformal loss function CE loss by adding a new term that quantifies uncertainty via conformal prediction. CALS [27] exploits the Augmented Lagrangian Multiplier method to adaptively learn the weights of the penalties to balance each term in the new loss function. Separately, class uncertainty-aware (CUA) loss [19] tailored object detection introduces the uncertainty of each class to augment the loss value when prediction results are uncertain. Further, Ding et al. [7] developed an uncertainty-aware loss for selective medical image segmentation that considers uncertainty in the training process to directly maximize the accuracy on the confident segmentation subset, rather than the accuracy on the whole dataset. Differently from the aforementioned works, our approach takes a distinct path by focusing on leveraging a pre-trained model to deliberately increase the uncertainty associated with incorrectly classified inputs without degrading the overall error rate.
Further, through our novel method, the objective is to empower Machine Learning (ML) systems to recognize possible model misclassification in production and take customized actions accordingly. This idea can be further extended to adversarial training scenarios, where the deliberate increase in uncertainty for misclassified adversarial inputs enhances the system’s ability to detect and respond to potential attacks in production.
@@ 3 Error-Driven Uncertainty Aware Training
DNNs often exhibit overconfidence in their predictions [11, 2]. Thus, in this work, we address this problem by proposing a loss function called Error-Driven Uncertainty Aware Training that shifts the focus toward refining the uncertainty guarantees of a pre-trained model by increasing the uncertainty associated with misclassifications while reducing the error rate and uncertainty of correct predictions. We assert that if a model reaches a stagnation phase in terms of quality and the pursuit of further training offers negligible benefit, it should, at the very least, produce high uncertainty values for incorrect predictions.
In order to achieve our design goals, we start the process by querying a pre-trained model to determine which inputs of the training dataset are wrong and correctly classified. Subsequently, we create two sets, one with incorrect classified inputs $\mathcal{W}$ and the other with the correct ones $\mathcal{C}$. We resize the correct classified inputs’ set by randomly selecting samples until the two sets have the same size. Then, in order to reduce overfitting, we mix wrong and correct classified inputs. Our approach employs distinct loss functions for each set. Since our objective is to deliberately increase the uncertainty of misclassifications, we minimize the CE and maximize the uncertainty (determined via MC dropout and PE) for the wrong-classified inputs while, for the correct-classified inputs, we minimize the CE and the uncertainty, i.e.,
$L_{\text{EUAT}}(f_{\theta}(\textbf{x}),\textbf{y})\text{=}\begin{cases}L_{\text{CE}}(f_{\theta}(\textbf{x}),\textbf{y})\text{-}L_{U}(f_{\theta}(\textbf{x}),\textbf{y})&\forall(x,y)\in\mathcal{W}\\ L_{\text{CE}}(f_{\theta}(\textbf{x}),\textbf{y})\text{+}L_{U}(f_{\theta}(\textbf{x}),\textbf{y})\hskip 5.69054pt&\forall(x,y)\in\mathcal{C}\end{cases}$ (1)
@@
$L_{\text{CE}}(f_{\theta}(x),y)=-\frac{1}{K}\sum_{i=1}^{K}t(x_{i})\log(f_{\theta}(x_{i})),$ (2)
($t(x)$ denotes the true label given the input $x$), and the uncertainty loss is measured by resorting to predictive entropy $H$
$L_{U}(f_{\theta}(x),y)=H[P(y|x)]=-\sum_{y\in\mathcal{Y}}P(y|x)\log P(y|x),$ (3)
where $P(y|x)$ is the model’s output distribution over the set of possible outcomes $\mathcal{Y}$ obtained via MC dropout by approximating the model’s output predictions using the average across parameters $\theta_{i}$ sampled from a dropout distribution
$p(y|\mathcal{D},x)\approx\frac{1}{N}\sum_{i=1}^{N}p(y|\theta_{i},x).$ (4)
We decided to also minimize uncertainty for correct-classified inputs in order to avoid a peak in the global model’s uncertainty. By using this approach, the model is able to better separate the erroneous and correct predictions using the uncertainty of each individual forecast. We also implement an early stopping policy based on the different evaluation metrics considered. It should also be noted that this proposed loss function is differentiable and continuous. The pseudo-code of the method proposed is described in Algorithm 1.
@@
1:  Input: model $f$, training set $\mathcal{S}$, optimizer opt
2:  Output: model $f$
3:  no_inputs = 0
4:  while True do $\triangleright$ spot condition
5:     $\mathcal{W}$, $\mathcal{C}$leftarrow $f$($\mathcal{S}$) $\triangleright$ Wrong $\mathcal{W}$ correct $\mathcal{C}$ classified sets
6:     $\mathcal{C}$leftarrow Random($\mathcal{C}$, $|\mathcal{W}|$)
@@
9:     $L(f(\textbf{x}),\textbf{y})$=$L_{\mathcal{W}}(f(\textbf{x}),\textbf{y})$+$L_{\mathcal{C}}(f(\textbf{x}),\textbf{y})$triangleright Add losses
10:     $L$.backward() $\triangleright$ Gradient estimation
11:     opt.step() $\triangleright$ Back-propagation computation
12:     no_inputs += ($|\mathcal{W}|$ + $|\mathcal{C}|$)
13:     if no_inputs $>$ epoch_size  then
14:        $f$.test() $\triangleright$ Test $f$ if trained with the same amount of data as the dataset
15:     end if
16:  end while
17:  return $f$
By resorting to EUAT, which effectively distinguishes between the uncertainty of accurate and erroneous predictions, we enhance the quality of uncertainty estimates and ultimately improve the reliability and trustworthiness of the model. Operationalizing the model in production involves assessing the uncertainty associated with each prediction. When the prediction falls below an uncertainty threshold, the model outcome can be trusted. Otherwise, when the uncertainty is above the threshold, the prediction is untrustworthy, warranting further scrutiny or review by human evaluators. Consequently, tuning the uncertainty threshold becomes fundamental. Akin to the implementation of the early stopping policy, we also explore various evaluation metrics to select the threshold.
In this work, we have considered classification tasks using different image recognition models. Thus, given the model’s prediction, we can promptly categorize it as either accurate or erroneous, and place it in the respective set. The early stopping policy and threshold selection are individually optimized for each metric.
We also consider a binary classification problem where the high-uncertainty output predictions that are likely to be wrongly classified can be inverted to the opposite class. Given the particularities of this problem, we tune the uncertainty threshold for EUAT differently by maximizing the number of correct predictions (i.e., true positives and true negatives) that are below the threshold and the number of incorrect predictions that are above the threshold (false positives and false negatives). Then, in production, when the outputted uncertainty is larger than the threshold, we can directly flip the prediction for the opposite class, and this way improve the model’s quality.
At last, we can also use EUAT to perform adversarial training. It should be noted that our approach does not aim to identify adversarial attacks using uncertainty; rather, when applied to AT focuses on identifying misclassifications based on the output uncertainty, regardless of whether they result from an attack or not. Equivalently, we can separate the wrong and correct predictions (clean and adversarial) in the two sets and train directly with our loss function. Since the function is differentiable, we can solve the optimization problem to find the perturbation using adversarial training methods like Fast Gradient Descent Method (FGSM) [13] or Project Gradient Descent (PGD) [30]. Similar to Smith and Gal [42], we acknowledge that neither dropout nor our method alone can be considered a reliable adversarial defense. Rather, we advocate for the joint use of all these methods when training a model adversarially in order to achieve better guarantees and trade-offs of accuracy, robustness, and uncertainty, yielding resilient models harder to attack.
@@ 4.1 Experimental Setup, Benchmarks, and Baselines
To evaluate EUAT, we employed four models and datasets widely used in the image recognition domain namely, ResNet50 [16] with ImageNet [40], Wide-ResNet-28x10 [46] with Cifar100 [21], ResNet18 with Cifar10, and ResNet18 with SVHN [35]. We also considered a binary classification model (using ResNet18 with Cifar10 to verify if there is a cat in an image), and an out-of-distribution (OOD) detection task, where corrupted inputs using distributional data shifts are used to evaluate the model. Finally, we evaluate our approach in adversarial training settings using three models/datasets mentioned above in the image recognition domain.
All the models and training procedures were implemented in Python3 via the Pytorch framework and trained using a single Nvidia RTX A4000.
To train the models, we used a dropout rate of 0.3 and resorted to stochastic gradient descent to minimize the loss function using a momentum of 0.9 and a batch size of 64 for all the models, a learning rate of 0.01 and weight decay of $10^{-5}$ for ResNet50/ImageNet and 0.1 and 0 for the remaining ones, respectively. Before training the model using EUAT, we pre-trained the models using CE loss and then decreased the learning rate by $10^{3}\times$ when applying EUAT. Additionally, we exploited automatic mixed precision to train the ResNet50/ImageNet and Wide-ResNet/Cifar100. The models were trained during 60 epochs (except in the binary classification problem where it was trained for 200 epochs). More in detail, we pre-trained the models for 30 epochs (100 epochs in the binary classification problem) before starting the second phase of training where we applied the EUAT. The implementation of the training pipeline and additional information to ensure the reproducibility of results are provided in the supplemental material.
We compared EUAT against the CE loss, model calibration, DEUP [23], an ensemble of five learners [24], CALS [27], and a loss function incorporating both CE and PE (CE+PE) [41]. To calibrate the model and train DEUP’s additional error predictor, we created a validation set comprising 10% randomly selected samples from the test set. Further, we resort to Isotonic regression [45] to calibrate the model, achieving superior results compared to other methods like Platt scaling [39], temperature scaling [14], and beta calibration [22]. Although we experimented DEUP with different validation set sizes, to ensure fairness, we maintained consistency by employing the same validation set size in both cases.
@@
We also evaluate the models using the correlation between the residuals of the model and predicted uncertainties (Corr. w/ res.) [23], and the Wasserstein distance of the uncertainty between the sets of correct and wrong predictions (Wasser. dist.). At last, we report the ECE and the model’s error. The uncertainty/confidence of the models is always computed via MC dropout using the normalized PE, except for DEUP, which resorts to the loss values of the base model to estimate the quality of its predictions (for a fair comparison, after testing the model trained with DEUP, we had to normalize the loss values). Moreover, we optimize each metric considered independently by employing an early stopping policy and tuning the uncertainty threshold, and we report the best values obtained for each metric.
@@
Standard Settings Adversarial Settings

### v2

@@ Abstract
Neural networks are often overconfident about their predictions, which undermines their reliability and trustworthiness. In this work, we present a novel technique, named Error-Driven Uncertainty Aware Training (EUAT), which aims to enhance the ability of neural classifiers to estimate their uncertainty correctly, namely to be highly uncertain when they output inaccurate predictions and low uncertain when their output is accurate. The EUAT approach operates during the model’s training phase by selectively employing two loss functions depending on whether the training examples are correctly or incorrectly predicted by the model. This allows for pursuing the twofold goal of i) minimizing model uncertainty for correctly predicted inputs and ii) maximizing uncertainty for mispredicted inputs, while preserving the model’s misprediction rate. We evaluate EUAT using diverse neural models and datasets in the image recognition domains considering both non-adversarial and adversarial settings. The results show that EUAT outperforms existing approaches for uncertainty estimation (including other uncertainty-aware training techniques, calibration, ensembles, and DEUP) by providing uncertainty estimates that not only have higher quality when evaluated via statistical metrics (e.g., correlation with residuals) but also when employed to build binary classifiers that decide whether the model’s output can be trusted or not and under distributional data shifts.
@@ 2 Related Work
One important foundation of these works lies in the computation of uncertainty. Uncertainty in DNNs plays a key role in quantifying the reliability and robustness of their predictions. There are two main types of uncertainty: epistemic uncertainty, associated with the model’s lack of knowledge or data, and aleatoric uncertainty, linked to the inherent randomness and unpredictability within the training data [43]. Various metrics are employed to measure these types of uncertainty such as predictive entropy (PE) and mutual information (MI) [42, 25]. However, quantifying uncertainty with DNNs is a challenging task. Bayesian methods [34, 29] can directly provide an estimate of the uncertainty by parameterizing the parameters of the network with distributions. However, training BNNs usually comes with a prohibitive cost. Thus, several approximations have been developed. Monte Carlo (MC) dropout [10], which is one of the most popular techniques for uncertainty quantification on DNNs [1, 4], adopts a Bayesian approximation to estimate the uncertainty by sampling multiple dropout masks and aggregating the predictions. Various statistical metrics of the aggregated predictions (such as PE, MI, and variance [5, 9, 26]) can then be used to quantify models’ uncertainty. Additionally, Markov Chain Monte Carlo (MCMC) methods [12] offer another avenue for drawing the posterior distribution (albeit, those require a long time to converge to the final distribution [1, 34]). Further, Variational Inference (VI) is a technique used to estimate the uncertainty of BNNs by approximating the posterior distribution over the model’s weights, which are treated as random variables with associated probability distributions. Training the network to approximate these distributions enables the capture of inherent uncertainty in the model’s predictions. Moreover, one can also resort to models that can directly output uncertainty estimations (such as Deep Gaussian Processes [6]), or use Laplace approximations for uncertainty  [30], or deep ensemble methods that offer yet another approach to estimate model uncertainty [24, 1], by aggregating the outputs of each learner in the ensemble and computing relevant metrics such as the entropy or MI. At last, DEUP [23] trains a new simple estimator to predict the uncertainty of the base model exploiting its error/loss, and DUN [2] leverages the outputs of different layers of a DNN to measure the uncertainty.
@@
Several works investigated how to calibrate a pre-trained model by applying a post-processing phase to optimize model calibration, i.e., to align the output probabilities with the true likelihood of the predicted outcomes of events occurring [14]. Several techniques, such as Platt Scaling [39], Isotonic Regression [47], Temperature Scaling [14], or Beta Calibration [22], can be employed to fine-tune the probabilities outputted by the model ensuring a more accurate alignment with true outcome frequencies. Krishnan et al. [20] introduced an accuracy versus uncertainty calibration (AvUC) loss function devised to obtain well-calibrated uncertainties while simultaneously preserving or enhancing model accuracy, and they extended their approach by proposing a post-hoc calibration phase that combines Temperature Scaling and AvUC. In addition, Karandikar et al. [18] proposed an extension of ECE and AvUC by developing a soft version of the binning operation underlying these calibration-error estimators, and also extended their approach for post-processing calibration by optimizing the temperature parameter in the temperature scaling method based on the soft calibration error. Complementary, Gupta et al. [15] presented a binning-free calibration approach. However, these calibration methods have been shown to be sensitive to both the model and the validation set [28] and perform sub-optimally when faced with shifts in data distribution [37].
@@
One fundamental aspect of training DNNs is the choice of a loss function. Although Cross Entropy (CE) stands out as a common choice for addressing classification problems, it tends to increase the over-confidence of the resulting model [14, 32]. Thus, since accounting for predictive uncertainty during training improves model calibration [20], several loss functions, such as focal loss [27] or label smoothing [27], have been developed that extend existing ones by incorporating additional terms addressing model uncertainty. Similarly, Shamsi et al. [41] proposed two loss functions that extend the CE by adding a new term to address the model’s uncertainty that can be determined through the PE or the ECE. Einbinder et al [8] introduced an uncertainty-aware conformal loss function by adding a new term that quantifies uncertainty via conformal prediction. CALS [28] exploits the Augmented Lagrangian Multiplier method to adaptively learn the weights of the penalties to balance each term in the new loss function. Separately, class uncertainty-aware (CUA) loss [19] tailored object detection introduces the uncertainty of each class to augment the loss value when prediction results are uncertain. Ding et al. [7] developed an uncertainty-aware loss for selective medical image segmentation that considers uncertainty in the training process to directly maximize the accuracy on the confident segmentation subset, rather than the accuracy on the whole dataset. Differently from the aforementioned works, our approach takes a distinct path by focusing on leveraging a pre-trained classification model to deliberately increase the uncertainty associated with incorrectly classified inputs without degrading the overall error rate.
Further, our novel method aims to empower Machine Learning (ML) systems to recognize possible model misclassification in production and take customized actions accordingly. This idea can be further extended to adversarial training scenarios, where the deliberate increase in uncertainty for misclassified adversarial inputs enhances the system’s ability to detect and respond to potential attacks in production.
@@ 3 Error-Driven Uncertainty Aware Training
This section introduces a new loss function, which we named Error-Driven Uncertainty Aware Training (EUAT), which aims to enhance the uncertainty estimation capability of a pre-trained model by leveraging a key principle: increasing the uncertainty associated with misclassifications while reducing the uncertainty of correct predictions and error rate of both correctly and incorrectly classified inputs.
The pseudo-code of EUAT is shown in Algorithm 1. We start the process by querying a pre-trained model to determine which inputs of its training set are wrongly and correctly classified (see lines 3 and  4). This leads to creating two sets, one containing the incorrect classified inputs $\mathcal{W}$ and the other containing the correct ones $\mathcal{C}$. We use stratified sampling on the target class in order to produce two subsets of equal size. Since we assume to use EUAT on pre-trained models, where typically the error rate on the training set is lower than 50%, i.e., $|\mathcal{C}|>|\mathcal{W}|$, this procedure typically entails sampling the set of correctly classified inputs ($\mathcal{C}$) in order to reduce its size to match the size of the set of mispredictions ($\mathcal{W}$). Then, in order to reduce overfitting, we mix wrong and correct classified inputs and ensure that the batches used for training with EUAT contain a balanced number of samples from $\mathcal{C}$ and $\mathcal{W}$.
EUAT employs distinct loss functions for each set. Since our objective is to deliberately increase the uncertainty of misclassifications, we minimize the CE and maximize the uncertainty for the wrong-classified inputs while, for the correct-classified inputs, we minimize the CE and the uncertainty, i.e.,
$L_{\text{EUAT}}(f_{\theta}(\textbf{x}),\textbf{y})\text{=}\begin{cases}L_{\text{CE}}(f_{\theta}(\textbf{x}),\textbf{y})\text{-}L_{U}(f_{\theta}(\textbf{x}),\textbf{y})&\forall\langle\textbf{x},y\rangle\in\mathcal{W}\\ L_{\text{CE}}(f_{\theta}(\textbf{x}),\textbf{y})\text{+}L_{U}(f_{\theta}(\textbf{x}),\textbf{y})\hskip 5.69054pt&\forall\langle\textbf{x},y\rangle\in\mathcal{C}\end{cases}$ (1)
@@
$L_{\text{CE}}(f_{\theta}(\textbf{x}),y)=-\frac{1}{K}\sum_{i=1}^{K}t(x_{i})\log(f_{\theta}(x_{i})),$ (2)
($t(x)$ denotes the true label given the input $x$ and $K$ represents the number of dimensions of x), and the uncertainty loss is measured by resorting to predictive entropy $H$
$L_{U}(f_{\theta}(\textbf{x}),y)=H[P(y|\textbf{x)}]=-\sum_{y\in\mathcal{Y}}P(y|\textbf{x})\log P(y|\textbf{x}),$ (3)
where $P(y|\textbf{x})$ is the model’s output distribution over the set of possible outcomes $\mathcal{Y}$ obtained via MC dropout, namely by approximating the model’s output predictions using the average across parameters $\theta_{i}$ sampled from a dropout distribution
$p(y|\mathcal{D},\textbf{x})\approx\frac{1}{N}\sum_{i=1}^{N}p(y|\theta_{i},\textbf{x}).$ (4)
We resort to MC dropout, since, as mentioned in Section 2, this approach has been shown to yield a more accurate estimation of the model’s uncertainty [10].
EUAT also includes a term that aims to minimize uncertainty for correctly classified inputs, which, intuitively, aims to instill confidence in the model in the regions of the input space where the model is already operating correctly. Further, by pushing the model’s uncertainty for correct and incorrect predictions in opposite directions, EUAT aims to ease the discrimination of erroneous and correct predictions via the uncertainty of the model’s forecasts. It should also be noted that this proposed loss function is differentiable and continuous, thus enabling its use with classical back-propagation-based training methods.
Note that at the end of each iteration of the while loop (line 2), the model’s weights are updated. This affects the model’s predictions and, consequently, also the sets $\mathcal{C}$ and $\mathcal{W}$, which is the reason why we update these two sets upon each iteration. Additionally, EUAT can be combined with commonly used stopping algorithms to determine when to stop the training process.
@@
1:  Input: model $f$, training set $\mathcal{S}$, optimizer opt , batch size $B$
2:  while StopCondition() $\neq$ True do
3:     $\mathcal{C}$ = $\langle x,y\rangle\in\mathcal{S}$ s.t. $y$=$f(x)$} $\triangleright$ Correct predictions set
4:     $\mathcal{W}$ = $\langle x,y\rangle\in\mathcal{S}$ s.t. $y\neq f(x)$} $\triangleright$ Mispredictions set
5:     $\mathcal{C}$leftarrow Select at random a subset of $\mathcal{C}$ of size |$\mathcal{W}$|
6:     for $i$ = 1, $\dots$, $\lceil$(|$\mathcal{W}$|+|$\mathcal{C}$|)/$B\rceil$ do
7:        $\triangleright$ $\mathcal{W}_{i}$ and $\mathcal{C}_{i}$ have size $B$/2, so the $i$-$t$h batch has size $B$ Let $\mathcal{W}_{i}$ and $\mathcal{C}_{i}$ be the $i$-$th$ mini-batch of $\mathcal{W}$ and $\mathcal{C}$, resp.
@@
10:        $L(f(\textbf{x}),\textbf{y})$=$L_{\mathcal{W}}(f(\textbf{x}),\textbf{y})$+$L_{\mathcal{C}}(f(\textbf{x}),\textbf{y})$triangleright Add losses
11:        $L$.backward() $\triangleright$ Compute and backpropagate the gradient
12:        opt.step() $\triangleright$ Gradient descent updating model’s parameters
13:     end for
14:  end while
15:  return $f$
@@ 4.1 Experimental Setup, Benchmarks, and Baselines
Datasets and Models.
In our experimental study, we employed four models and datasets widely used in the image recognition domain namely, ResNet50 [16] with ImageNet [40], Wide-ResNet-28x10 [48] with Cifar100 [21], ResNet18 with Cifar10, and ResNet18 with SVHN [35]. We also considered a binary classification model (using ResNet18 with Cifar10 to verify if there is a cat in an image), and an out-of-distribution (OOD) detection task, where corrupted inputs using distributional data shifts are used to evaluate the model. All these models make use of dropout layers and we configure them to use a dropout rate of 0.3. Finally, we evaluate our approach in adversarial training settings using the three models/datasets mentioned above in the image recognition domain.
Baselines.
We compared EUAT against the CE loss, model calibration, DEUP [23], an ensemble of five learners [24], CALS [28], and a loss function incorporating both CE and PE (CE+PE) [41]. To calibrate the model and train DEUP’s additional error predictor, we created a validation set comprising 10% randomly selected samples from the dataset. Further, we resort to Isotonic regression [47] to calibrate the model, which we have confirmed to achieve superior results when compared to other methods like Platt scaling [39], temperature scaling [14], and beta calibration [22]. Although we experimented DEUP with different validation set sizes, to ensure fairness, we maintained consistency by employing the same validation set size in both cases. For a fair comparison, we ensure each epoch has seen the same amount of data.
Evaluation metrics.
@@
We also evaluate the models using the correlation between the residuals of the model and predicted uncertainties (Corr. w/ res.) [23], and the Wasserstein distance [44] of the uncertainty distribution between the sets of correct and wrong predictions (Wasser. dist.). At last, we report the ECE and the model’s error. The uncertainty/confidence of the models is always computed via MC dropout using the normalized PE, except for DEUP, which resorts to the loss values of the base model to estimate the quality of its predictions (for a fair comparison, after testing the model trained with DEUP, we had to normalize the loss values).
Additional implementation details.
To train the models, independently of the considered solution, we use stochastic gradient descent to minimize the loss function using a momentum of 0.9 and a batch size of 64 for all the models, a learning rate of 0.01, and a weight decay of $10^{-5}$ for ResNet50/ImageNet and 0.1 and 0 for the remaining ones, respectively.
Before training the model using EUAT, we pre-trained the models using CE loss and then decreased the learning rate by $10^{3}\times$ when applying EUAT. Additionally, we exploited automatic mixed precision to train the ResNet50/ImageNet and Wide-ResNet/Cifar100. The models were trained during 60 epochs (except in the binary classification problem where it was trained for 200 epochs). More in detail, we pre-trained the models for 30 epochs (100 epochs in the binary classification problem) before starting the second phase of training where we applied EUAT (for all the remaining epochs). We use a validation set (corresponding to 10% of the dataset) in order to tune the uncertainty threshold that is used to discriminate the correctly and incorrectly classified inputs (this threshold is necessary, e.g., to compute the Uncertainty Confusion Matrix). Also, in EUAT, we save the model produced at the end of each training epoch, evaluate its performance on the validation set, and select the one that maximizes the considered performance metric. For fairness, also for the other baselines, we use the same validation-based method to optimize the tuning of the uncertainty threshold as well as the model’s selection during the training phase.
The implementation of the training pipeline and additional information to ensure the reproducibility of results are provided in the public repository [^1] https://github.com/pedrogbmendes/EUAT. All the models and training procedures were implemented in Python3 via the Pytorch framework and trained using a single Nvidia RTX A4000.
@@
Evaluation with clean data Evaluation with adversarial data